OpenStudy (anonymous):

Which values of x and y will satisfy y = 5x and 12x + 9y = 60

OpenStudy (anonymous):

y = 5x 12x + 9y = 60 Use substitution and solve the system 12x + 9(5x) = 60 12x + 45x = 60 57x = 60 x=20/19 y = 5(20/19) y= 100/19 @Conqueror right or nah ?

OpenStudy (conqueror):

Dunno

OpenStudy (anonymous):

@Kemist ?